A hyperbola has the equation
16x²-9y²+64x-90y =305
Find its center,foci,vertices, length of latus rectum and equation of the asymptotes
1
Expert's answer
2020-08-20T18:10:43-0400

"16x^2-9y^2+64x-90y=305\\
16(x^2+4x)-9(y^2+10y)=305\\
16(x^2+4x+4)-9(y^2+10y+25)=305+64-225"

"16(x+2)^2-9(y+5)^2=144\\
\frac{(x+2)^2}{9}-\frac{(y+5)^2}{16}=1"

h=-2 , k=5,

"a=\sqrt9=3\\
b=\sqrt{16}=4\\
c=\sqrt{a^2+b^2}=\sqrt{9+16}=\sqrt{25}=5"


Centre: (-2,-5),

vertices: (-5,5) and (1,-5),

foci: (-7,-5) and (3,-5),

asymptotes:

"y=-5\pm4\frac{(x+2)}{3}"


Length of latus rectrum ="\frac{2b^2}{a}\\"

"=\frac{2*16}{3}\\
=\frac{32}{3}\\
=10.66"